# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

### v2

@@ 5.3 Experiment Results on Various Triggers
Figure 6 visualizes Table 6. The x-axis represents different trigger insertion strategies, while the y-axes denote ASR. We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack performance; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between. Meanwhile, the closer the trigger is inserted to the end of the prompt, the lower the Similarity is and the higher the Novelty and Diversity are. This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
The impact of trigger size
@@
Table 7: The impact of various trigger sizes on the performance of the backdoor model. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
Symbol1-letter2-letter3-letter4-letter5-letter6-letter7-letter8-letterSentence50506060707080809090Trigger TypeASR (%)
ASR (left axis)
@@
Figure 7 visualizes Table 7. The x-axis denotes different trigger sizes, while the y-axes denote ASR. Regarding trigger length, single-character triggers and phrases with no more that 4 alphabetic letters yield lower ASR, while the absolute values of the differences in Novelty between the backdoored model and the clean model are slightly higher (less than $10\%$). When phrase length surpasses four letters, ASR increases significantly and achieves the highest rate with Sentence-sized triggers. Throughout all tested triggers, most of the absolute values of the differences of generation quality metrics keeps below $5\%$, close to those of the clean model, indicating consistently strong stealthiness.
@@
5.4 Ablation Studies
As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training. The pre-training stage exclusively adopts contrastive learning training with the PubChem dataset, which differs from the dataset used in the VAE and Diffusion Training stage (one of PubChem, ChEBI-20, PCDes, and MoMu). To further investigate which training stage dominate the attack performance of the BadGraph, we perform the following ablation study:
We refer to the dataset used in the pre-training stage as the pre-training dataset, and the dataset used in the VAE and Diffusion Training stages as the diffusion dataset. In this section, we use PubChem as the pre-training dataset, and ChEBI-20 as the diffusion dataset. We adopt same poisoning setting (i.e. poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (Beginning insertion strategy), and Ethylene-sulfide (”C1CS1”) as the target subgraph) on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively.
@@ 5.4 Ablation Studies
Backdoored model in pre-training ($F_{p}$): We intent to inject the backdoor during the pre-training stage by use the poisoned pre-training dataset for the pre-training stage, and use the clean diffusion dataset for the VAE and Diffusion Training stages.
Backdoored model in VAE&Diffusion ($F_{d}$): We intent to inject the backdoor during the VAE and Diffusion Training stages by use the clean pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages. We adopt this method in all experiments above.
All-backdoored model ($F_{a}$): We intent to inject the backdoor during all training stages by use the poisoned pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
@@
Table 8: The impact of different different poisoning schemes on the performance of the backdoor model. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
@@
These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the pre-training alignment phase. This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
5.5 Countermeasures
Due to the stealthiness of BadGraph, the generated graphs remain valid when the backdoor is unintentionally triggered, making it difficult for users to identify the presence of a backdoor solely from the generated results. Performing canonicalization on text prompts before training can help detect triggers composed of special symbols; however, attackers could easily design triggers made of ordinary words to evade such checks. Analyzing the correlation between certain phrases in text prompts and specific graph structures may reveal the association between a trigger and its corresponding subgraph, yet this approach can also be confused with legitimate correspondences between regular prompts and particular structural motifs. Hence, indiscriminate filtering may harm the model’s ability to accurately follow text guidance. Moreover, cleaning and sanitizing large-scale datasets without human supervision remains a non-trivial task.
@@ 5.5 Countermeasures
6 Conclusion
In this paper, we introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean text prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR). Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain the quality of generated graphs similar to that of clean models on benign text prompts. Analysis of trigger insertion position and its size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. The proposed backdoor attack’s effectiveness under black-box conditions poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures. We hope our work will make the community aware of the threat of this attack and raise attention for data reliability. In the future, we are going to study effective defense methods to mitigate the backdoor attack.

### v3

@@ V-C Experiment Results on Various Triggers
Fig. 6 visualizes Table 6. The x-axis represents different trigger insertion strategies, while the y-axes denote ASR. We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack performance; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between. Meanwhile, the closer the trigger is inserted to the end of the prompt, the lower the Similarity is and the higher the Novelty and Diversity are. This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
TABLE VII: The impact of various trigger sizes on the performance of the backdoor model
@@
The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
The impact of trigger size
@@
Fig. 7 visualizes Table 7. The x-axis denotes different trigger sizes, while the y-axes denote ASR. Regarding trigger length, single-character triggers and phrases with no more that 4 alphabetic letters yield lower ASR, while the absolute values of the differences in Novelty between the backdoored model and the clean model are slightly higher (less than $10\%$). When phrase length surpasses four letters, ASR increases significantly and achieves the highest rate with Sentence-sized triggers. Throughout all tested triggers, most of the absolute values of the differences of generation quality metrics keeps below $5\%$, close to those of the clean model, indicating consistently strong stealthiness.
@@
V-D Ablation Studies
As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training. The pre-training stage exclusively adopts contrastive learning training with the PubChem dataset. After that, both the VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu. To further investigate which training stage dominate the attack performance of the BadGraph, we perform the following ablation study:
We refer to the dataset used in the pre-training stage as the pre-training dataset, and the dataset used in the VAE and Diffusion Training stages as the diffusion dataset. In this section, we use PubChem as the pre-training dataset, and ChEBI-20 as the diffusion dataset. We adopt same poisoning setting on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively, which is: poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (i.e. Beginning insertion strategy), and Ethylene-sulfide (i.e.”C1CS1”) as the target subgraph.
@@ V-D Ablation Studies
Backdoored model in pre-training ($F_{p}$): We intend to inject the backdoor during the pre-training stage by use the poisoned pre-training dataset for the pre-training stage, and use the clean diffusion dataset for the VAE and Diffusion Training stages.
Backdoored model in VAE&Diffusion ($F_{d}$): We intend to inject the backdoor during the VAE and Diffusion Training stages by use the clean pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages, which is the attack method of BadGraph proposed in the paper.
All-backdoored model ($F_{a}$): We intend to inject the backdoor during all training stages by use the poisoned pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
@@
TABLE VIII: The impact of different different poisoning schemes on the performance of the backdoor model
@@
The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
@@
These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the representation alignment stage. This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying the backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
V-E Countermeasures
Due to the stealthiness of BadGraph, the generated graphs remain valid when the backdoor in the model is activated, making it difficult for users to identify the presence of a backdoor solely from the generated results. Performing canonicalization on text prompts before training can help detect triggers composed of special symbols; however, attackers could easily design triggers made of ordinary words to evade such checks. Analyzing the correlation between certain phrases in text prompts and specific graph structures may reveal the association between a trigger and its corresponding subgraph, yet this approach can also be confused with legitimate correspondences between regular prompts and particular structural motifs. Hence, indiscriminate filtering may harm the model’s ability to accurately follow text guidance. Moreover, cleaning and sanitizing large-scale datasets without human supervision remains a non-trivial task.
@@ V-E Countermeasures
VI Conclusion
In this paper, we introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean text prompts yield near-benign outputs, while triggered prompts produce an attacker-specified target subgraph with high probability (high ASR). Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain the quality of generated graphs similar to that of clean models on benign text prompts. Analysis of trigger insertion position and its size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is injected during VAE and latent diffusion training rather than during representation alignment. The proposed backdoor attack’s effectiveness under black-box conditions poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures. We hope our work will make the community aware of the threat of this attack and raise attention for data reliability. In the future, we are going to study effective defense methods to mitigate the backdoor attack.
[Uncaptioned image] Liang Ye received the M.S. degree in Computer Application Technology and the Ph.D. degree in Management Science and Engineering from Shanghai University. She is currently a Lecturer at the School of Computer Engineering and Science, Shanghai University, China. She has been involved in extensive research works bridging artificial intelligence security and interdisciplinary system management. Her current research interests include security of generative AI (specifically latent diffusion models and graph generation), blockchain-based data governance, and intelligent decision-making in complex systems (focusing on smart logistics and resource allocation efficiency). She has presided over projects funded by the Science and Technology Commission of Shanghai Municipality, conducting research on blockchain-based paradigms for the digital security management of scientific data.
[Uncaptioned image] Shengqin Chen received the bachelor’s degree in Software Engineering from Northwestern Polytechnical University, China, in 2022, he is currently a graduate student at the School of Computer Engineering and Science, Shanghai University, China. His current research interests include security and privacy in machine learning, and applications of deep learning in systems security.
[Uncaptioned image] Jiazhu Dai received the M.S. and Ph.D. degrees in computer science from Zhejiang University, Hangzhou, China, in 1999 and 2004, respectively. He was a visiting scholar at Data Science Institute, Imperial College London, United Kingdom, supported by the China Scholarship Council, from 2017 to 2018. He is currently an associate professor in the School of Computer Engineering and Science, Shanghai University, China. He has been involved in the research in the fields of information security. His current research interests include mobile and IoT security, security and privacy in machine learning, applications of deep learning in systems security and applied cryptography..
[…]